OpenStudy (anonymous):

A large totem pole near Kalama, Washington, is 106 feet tall. At a particular time of day, the angle of elevation to the top of the pole from the end of the shadow it casts is 23 degrees. What is the length of the shadow?

OpenStudy (anonymous):

 take the tan function of 23, so tan(23) = 106/x bring the x over so xtan(23) = 106 bring the tan over to get x alone since that's what you're finding so x = 106/tan(23) plug that into a calculator and that's the length of the shadow